# Cross-Modal and Uncertainty-Aware Agglomeration for Open-Vocabulary 3D Scene Understanding

### v1

@@ Abstract
The lack of a large-scale 3D-text corpus has led recent works to distill open-vocabulary knowledge from vision-language models (VLMs). However, these methods typically rely on a single VLM to align the feature spaces of 3D models within a common language space, which limits the potential of 3D models to leverage the diverse spatial and semantic capabilities encapsulated in various foundation models. In this paper, we propose Cross-modal and Uncertainty-aware Agglomeration for Open-vocabulary 3D Scene Understanding dubbed CUA-O3D, the first model to integrate multiple foundation models—such as CLIP, Dinov2, and Stable Diffusion—into 3D scene understanding. We further introduce a deterministic uncertainty estimation to adaptively distill and harmonize the heterogeneous 2D feature embeddings from these models. Our method addresses two key challenges: (1) incorporating semantic priors from VLMs alongside the geometric knowledge of spatially-aware vision foundation models, and (2) using a novel deterministic uncertainty estimation to capture model-specific uncertainties across diverse semantic and geometric sensitivities, helping to reconcile heterogeneous representations during training. Extensive experiments on ScanNetV2 and Matterport3D demonstrate that our method not only advances open-vocabulary segmentation but also achieves robust cross-domain alignment and competitive spatial perception capabilities. The code will be available at https://github.com/TyroneLi/CUA_O3D.
@@ 1 Introduction
3D scene understanding serves as a crucial perception component for a wide array of real-world applications to help models better understand the physical world, including robot navigation, autonomous vehicles, and virtual reality [23, 82, 50, 5]. Typical approaches necessitate a dataset of semantically annotated point clouds, which is both time-consuming (e.g., 22.3 minutes for annotating a single scene with 20 classes [15]) and unable to encompass all possible existing categories, thereby limiting their practical utility. To address this limitation, recent works have explored the open-vocabulary 3D scene understanding setting [27, 35, 64, 17], aiming to localize and recognize arbitrary object classes. This objective is achieved by leveraging Vision-Language Models (VLMs) [66, 36] that, having been pre-trained on billions of image-text pairs [76], can gauge the alignment between textual and visual inputs, facilitating a range of 2D open-vocabulary tasks [25, 102, 6, 22, 41]. To adapt these models for 3D downstream tasks (e.g., semantic segmentation), various strategies, predominantly based on distilling multi-view 2D visual features into a 3D-specific model [27, 35, 64, 17, 87], have been employed.
@@
To explore various foundation model properties when fusing multi-view posed image features to 3D space and enabling model distillation, we first conduct a pilot study to analyze and compare the heterogeneous results shown in Fig. 1. We can observe that the distribution in terms of the feature embeddings from each 2D foundation model mainly follows a gaussian-like feature distribution. At the same time, when clustering these features into specific clusters, the results illustrate heterogeneous and complementary effects across different foundation models. This is also due to the inconsistency across different posed images when the 2D model encounters complex contexts. This motivates us to develop a new method to harmonize this heterogeneous knowledge into a 3D model and handle such noisy inconsistency from the fused 2D feature embeddings.
@@
In this paper, we present Cross-modal and Uncertainty-aware Agglomeration for Open-vocabulary 3D Scene Understanding, dubbed CUA-O3D, the first model to integrate multiple foundation models into one 3D model for scene understanding. Deterministic uncertainty estimation is further introduced to adaptively distill and harmonize the heterogeneous 2D feature embeddings from these models. We show that there are potential inconsistencies yet complementarity when attempting to distill different foundation models. Based on our pilot study, we first propose to leverage distillation loss to supervise 3D model training given several available 2D feature embeddings. Our 3D model consists of independent projection layers to be mapped with one VLM or visual foundation model under feature supervision which helps with reconciling the entanglements from heterogeneous distributions. To resolve the potential noises from 2D models which mainly come from insufficient contexts, a novel deterministic uncertainty estimation is tailored to adaptively weight the knowledge distillation which can be modeled like a gaussian likelihood that follows the distributions as shown at the top of Fig. 1. Specifically, regarding each projection layer to be mapped with the specific 2D model, we tailor an observation noise scalar prediction independently to capture how much the noise is contained in feature supervision during training, which is termed uncertainty-aware learning. Moving one step forward, as can be observed the distribution in terms of Stable Diffusion shows a bit shift away from the center scale and usually carries broader value ranges and heavy-tailed (“spike”) values in the projected feature embeddings. A de-mean operation is then adopted to re-center the feature scales, being able to reduce the impact from anomaly points while still allowing points with small scale to guide the 3D model training.
@@
  • We analyze the heterogeneous yet complementary feature embeddings from multiple 2D models and incorporate both semantic- and geometric-aware knowledge into a 3D model.
  • We further propose a deterministic uncertainty estimation to enable the 3D model output independent observation scalar to capture the noise and resolve the heterogeneity from various feature supervisions.
@@ 2 Related Works
Open-Vocabulary (OV) 3D scene understanding advances over the previous large corpus of close-set approaches [9, 53, 70, 3, 98, 52, 87], allowing robust zero-shot reasoning and alleviating the need for annotations. Recent advances in Visual-Language Models (VLMs) [66, 12, 2, 36] have driven OV models towards remarkable levels of robustness with numerous emerging approaches tackling OV in image semantic segmentation [22, 6, 51, 86], object detection [104, 4], and recently universal segmentation [93]. Differently, OV for 3D scene understanding (OV3D) is limited in the data availability for training a purely fundamental 3D VLM. Alternatively, the community achieves OV3D by distilling zero-shot knowledge from recent VLMs [66, 6] and by mapping point cloud features to a queryable CLIP space. In 3D semantic segmentation, ConceptFusion [35] fuses VLM representations from multiple views into 3D points. Some methods [79] extends OV to 3D instance semantic segmentation based on CLIP [66] or SAM [40] to align the 3D space with language space while forcing instance-mask constraint. However, the recent OV3D methods heavily rely on the zero-shot knowledge of the 2D VLMs without investigating the reliability of the projected 2D feature representation. In this work, we move forward and explore how to aggregate knowledge from various 2D foundation models.
@@
Knowledge distillation (KD) [29, 62] aims at training compact student models with the supervision from more powerful and larger teacher models [47]. Introduced by the first work [29], the student model is trained to mimic the prediction behavior of the teacher model and has been extensively explored in subsequent works [69, 91, 1, 97, 49, 99], which has been applied successfully in a wide range of tasks going from supervised-training [56, 85, 81], to network compression [75, 8, 65] and to domain adaptation [100, 30, 28]. Recently, following the explosion of Visual-Language Models like CLIP [66] and ALIGN [36], KD has been introduced to efficiently transfer knowledge between different modalities [47, 54, 25], and recently to bridge the gap between the text and 3D point cloud modalities [10, 90, 34, 96, 64, 17, 88]. Very recently, AM-RADIO [67] describes a general methodology for distilling multiple distinct foundation models into one, but still focus on only 2D domain. We primarily focus on studying and tackling the ambiguity of the distilled representations between image and point cloud modalities.
@@
Uncertainty estimation has been widely investigated in various tasks [37, 31, 42, 44, 60, 26, 57, 38] which is capable of addressing the problem of quantifying the uncertainty of predictions by model. Uncertainty Estimation can be broadly classified into aleatoric estimation [37, 83, 59] that usually dues to the underlying uncertainty in the measurement which utilizes the extra network to be trained from scratch to approximate a heteroscedastic distribution by maximizing the likelihood of the system, and epistemic estimation [24, 42, 21] that induces the uncertainty by the model parameters in low-data regimes as parameter estimation becomes noisy, respectively. In 3D point clouds, uncertainty estimation finds applications in incremental learning [89] and semantic segmentation [14], and domain adaptation [73, 84, 74]. Unlike previous works, we seek a deterministic uncertainty estimation for supervising 3D model training under uncertainty awareness for the ambiguity between image and point modalities.
@@ 3 Methodology
In this section, we first describe the preliminary open-vocabulary scene understanding task in Sec. 3.1. Then we elucidate inconsistent results across multi-view posed images to demonstrate the necessity of uncertainty-aware training to alleviate such issues and embrace various 2D foundation models in Sec. 3.2. Our Cross-Modal and Uncertainty-Aware Agglomeration (CUA-O3D) method will be depicted in Sec. 3.3.
@@ 3.1 Open-Vocabulary 3D Scene Understanding
In standard 3D semantic segmentation, the training set $\mathcal{T}_{train}$ includes point clouds and dense point-level annotations. Each point cloud $\mathcal{X}={p_{i}\in\mathbb{R}^{3},i\in[0,N-1]}$, consists of $N$ points $p_{i}$ with corresponding point-level annotations $\mathcal{Y}$. These annotations are assigned from a predefined set of class indices $\mathcal{K}=[1,\dots,K]$, where each index corresponds to a specific class name in the vocabulary $\mathcal{V}=[v_{1},\dots,v_{k}]$. Given $\mathcal{T}_{train}$ and $\mathcal{K}$, the objective is to learn a deep neural network correctly assigning a label from $\mathcal{K}$ to each point $p_{i}\in\mathcal{X}$. In contrast, open-vocabulary 3D semantic segmentation (OV3D) aims to segment $\mathcal{X}$ using an arbitrary vocabulary $\mathcal{V}$. Recent approaches achieve this with a pre-trained VLM [64, 79]. The VLM provides this common embedding space through two distinct encoders, namely, a vision encoder $f^{2D}:\mathcal{I}\rightarrow\mathbf{Z}$ and a text encoder $f^{txt}:\mathcal{V}\rightarrow\mathbf{Z}$. The common practice is to train a 3D vision encoder $\theta_{3D}$ to align to $f^{2D}$ embeddings, bridging the modality gap and defining $\theta_{3D}:\mathcal{X}\rightarrow\mathbf{Z}$. After training, $f^{txt}$ encodes $\mathcal{V}$, and class predictions are computed via similarity matching between point and vocabulary embeddings.
@@ 3.2 Preliminary Observation
We first conduct a simple qualitative study to analyze how embedding ambiguity affects $f^{2D}$ predictions before and after projection to the point cloud space based on commonly used Lseg [6]. We analyze the consistency of the predictions for the same object appearing in multi-view images. Given a pre-trained vision-language model $f^{2D}$ and multi-view images $\mathcal{I}$, we query $f^{2D}$ with known vocabularies $\mathcal{V}$ and report the qualitative results over multiple ScanNetV2 [15] view-images in Fig. 2. We notice that predictions are consistent for the classes wall, floor, and door while showing inconsistency over the class cabinet. After projection to the point cloud space (left), the projected prediction inherits this ambiguity, resulting in detrimental 3D model distillation. Regarding the incorporation of various 2D foundation models shown in Fig. 1, how to harmonize the heterogeneous characteristics also matters. This simple study highlights the need for a reliable uncertainty measure capturing embedding ambiguity that we hope to shed new light on the future works on this task.
@@ 3.3 Cross-Modal Agglomeration
Distillation agglomeration. The distillation phase aims to align the point embeddings $F^{3D}$ to image embeddings $F^{2D}$ obtained from each frozen pre-trained visual encoder $f^{2D}_{i}$. This is a common practice in recent OV approaches [79], which results in a shared embedding space between image, text, and point cloud modalities. Given a point cloud $\mathcal{X}$ paired with multi-view posed images $\mathcal{I}$, we learn 3D sparse features $F^{3D}$ by employing the well-established MinkowskiNet [13] as a 3D sparse convolutional encoder $\theta_{3D}$. The encoder $\theta_{3D}$ outputs a sparse set of point-wise feature vectors $F^{3D}=\theta_{3D}(\mathcal{X})$, where each feature is associated with an input point $p_{i}\in\mathcal{X}$. Similarly, we distill the multi-view image features from $\mathcal{I}$ using each frozen vision encoder $f^{2D}_{i}$ separately, where $\{f^{2D}_{i}\in f^{2D}|f^{2D}_{Lseg},f^{2D}_{DINOv2},f^{2D}_{StableDiffusion}\}$. The output of $f^{2D}$ is a set of feature vectors $F^{2D}_{i}=f^{2D}_{i}(\mathcal{I})$ where each feature vector in $F^{2D}$ is associated to an input pixel $u$. After projection in the homogeneous coordinates space, we redefine $F^{2D}_{i}$ as a set of point-wise image features, corresponding to each $f^{2D}_{i}$. We define the matches between each point $p$ and pixel $u$ by using the corresponding homogeneous coordinates $\tilde{p}$ and $\tilde{u}$, respectively. Once matches are established, we enforce the alignment between the predicted $F^{3D}_{i}$ and $F^{2D}_{i}$. Following [64] and training $\theta_{3D}$ to minimize the distillation loss $\mathcal{L}_{distill}$, our independent distillation losses are defined as,
@@
which corresponds to minimizing the distance between each $F^{3D}_{i}$ and $F^{2D}_{i}$, leading to final distillation loss as,
@@
Deterministic uncertainty estimation. As we analyzed before different 2D foundation models encapsulate unique characteristics and one single 2D model induces inherent inconsistency from multi-view posed image which necessitates appropriate measures to tackle. We then propose a simple yet effective deterministic uncertainty-aware observation scalar prediction to quantify embedding ambiguity within each cross-modal distillation, that learns the adaptive weights of various 2D feature supervisions under the cross-modal training.
Specifically, we devise the 3D model $\theta_{3D}$ with three independent noise scalar predictions $\sigma_{i}$  w.r.t each 2D model $f^{2D}_{i}$. The output from the probabilistic model with weight $W$ being analogous to the regression task can be modeled as gaussian likelihood as:
$p(y|f^{W}(x))=N(f^{W}(x),\sigma^{2}),$ (6)
@@
$p(y_{1},...,y_{K}|f^{W}(x))=\prod\limits_{i=1}^{K}p(y_{i}|f^{W}(x)),$ (7)
where index $i$ corresponds to the mapping for each 2D model, $\{f^{2D}_{Lseg},f^{2D}_{DINOv2},f^{2D}_{StableDiffusion}\}$. Based on the maximum likelihood modeling and taking the regression task as an example, the log-likelihood of the model can be optimized,
$logp(y|f^{W}(x))\propto-\frac{1}{2\sigma^{2}}||y-f^{W}(x)||^{2}-\log\sigma,$ (8)
where $\sigma$ denotes the model’s observation noise scalar, being responsible for capturing the inherent uncertainty within 2D feature supervisions, which is mainly caused by heterogeneous and noisy 2D feature embeddings from various 2D models. Assuming that we have three outputs corresponding to Lseg, DINOv2, and Stable Diffusion, each following gaussian-like distribution, we then have:
$\begin{split}p(y_{1},y_{2},y_{3}|f^{W}(x))&=\prod\limits_{i=1}^{3}p(y_{i}|f^{W}(x))\\ &=\prod\limits_{i=1}^{3}N(y_{i};f^{W}(x),\sigma_{i}^{2}),\end{split}$ (9)
and then, we formulate our training objective $\mathcal{L}_{distill}(W,\sigma_{1},\sigma_{2},\sigma_{3})$ for multiply mappings:
$\vspace{-0.2cm}\begin{split}&\mathcal{L}_{distill}=-\log p(y_{1},y_{2},y_{3}|f^{W}(x))\\ &\propto\frac{1}{2\sigma_{1}^{2}}\mathcal{L}_{cos\_lseg}+\frac{1}{2\sigma_{2}^{2}}\mathcal{L}_{l1}+\frac{1}{2\sigma_{3}^{2}}\mathcal{L}_{cos\_sd}+\log\sigma_{1}\sigma_{2}\sigma_{3},\end{split}$ (10)
@@
To the best of our knowledge, this is the first work investigating deterministic uncertainty-aware modeling to agglomerate various 2D foundation models into one single 3D model in pursuit of the potential foundational 3D model. In the last term of Eq. 10  w.r.t each supervision from the 2D model plays the role of learning adaptive weightings during the training. When the corresponding $\sigma_{i}$ denoting the uncertain noise parameter for $f^{2D}_{i}$ increases, the weight of $\mathcal{L}_{i}$ will decreases, while each $\sigma_{i}$ will avoid increasing too much. We visualize each $\sigma_{i}$ during the training in supplementary. To further maintain the stable training, we add a small constant $1.0$ to each $\sigma_{i}$ to avoid the loss of becoming negative during training,
$\begin{split}\log\sigma_{i}\to\log(1.0+\sigma_{i}).\end{split}$ (11)

### v2

@@ Abstract
The lack of a large-scale 3D-text corpus has led recent works to distill open-vocabulary knowledge from vision-language models (VLMs). However, these methods typically rely on a single VLM to align the feature spaces of 3D models within a common language space, which limits the potential of 3D models to leverage the diverse spatial and semantic capabilities encapsulated in various foundation models. In this paper, we propose Cross-modal and Uncertainty-aware Agglomeration for Open-vocabulary 3D Scene Understanding dubbed CUA-O3D, the first model to integrate multiple foundation models—such as CLIP, DINOv2, and Stable Diffusion—into 3D scene understanding. We further introduce a deterministic uncertainty estimation to adaptively distill and harmonize the heterogeneous 2D feature embeddings from these models. Our method addresses two key challenges: (1) incorporating semantic priors from VLMs alongside the geometric knowledge of spatially-aware vision foundation models, and (2) using a novel deterministic uncertainty estimation to capture model-specific uncertainties across diverse semantic and geometric sensitivities, helping to reconcile heterogeneous representations during training. Extensive experiments on ScanNetV2 and Matterport3D demonstrate that our method not only advances open-vocabulary segmentation but also achieves robust cross-domain alignment and competitive spatial perception capabilities. Project webpage: CUA-O3D.
@@ 1 Introduction
3D scene understanding serves as a crucial perception component for a wide array of real-world applications to help models better understand the physical world, including robot navigation, autonomous vehicles, and virtual reality [21, 82, 48, 4]. Typical approaches necessitate a dataset of semantically annotated point clouds, which is both time-consuming (e.g., 22.3 minutes for annotating a single scene with 20 classes [13]) and unable to encompass all possible existing categories, thereby limiting their practical utility. To address this limitation, recent works have explored the open-vocabulary 3D scene understanding setting [33, 64, 15, 54, 55], aiming to localize and recognize arbitrary object classes. This objective is achieved by leveraging Vision-Language Models (VLMs) [66, 34] that, having been pre-trained on billions of image-text pairs [76], can gauge the alignment between textual and visual inputs, facilitating a range of 2D open-vocabulary tasks [23, 102, 5, 20, 39]. To adapt these models for 3D downstream tasks (e.g., semantic segmentation), various strategies, predominantly based on distilling multi-view 2D visual features into a 3D-specific model [25, 33, 64, 15, 87], have been employed.
@@
To explore various foundation model properties when fusing multi-view posed image features to 3D space and enabling model distillation, we first conduct a pilot study to analyze and compare the heterogeneous results shown in Fig. 1. We can observe that the distribution in terms of the feature embeddings from each 2D foundation model mainly follows a gaussian-like feature distribution. At the same time, when clustering these features into specific clusters, the results illustrate heterogeneous and complementary effects across different foundation models. This is also due to the inconsistency across different posed images when the 2D model encounters complex image contexts. This motivates us to develop a new method to harmonize this heterogeneous knowledge into a 3D model and handle such noisy inconsistency from the fused 2D feature embeddings.
@@
In this paper, we present Cross-modal and Uncertainty-aware Agglomeration for Open-vocabulary 3D Scene Understanding, dubbed CUA-O3D, the first method to integrate multiple 2D foundation models into one 3D model for scene understanding. Deterministic uncertainty estimation is further introduced to adaptively distill and harmonize the heterogeneous 2D feature embeddings from these models. We show that there are potential inconsistencies yet complementarity when attempting to distill different foundation models. Based on our pilot study, we first propose to leverage distillation loss to supervise 3D model training given several available 2D feature embeddings. Our 3D model consists of independent projection layers to be mapped with one VLM or visual foundation model under feature supervision which helps with reconciling the entanglements from heterogeneous distributions. To resolve the potential noises from 2D models which mainly come from insufficient contexts, a novel deterministic uncertainty estimation is tailored to adaptively weight the knowledge distillation which can be modeled like a gaussian likelihood that follows the distributions as shown at the top of Fig. 1. Specifically, regarding each projection layer to be mapped with the specific 2D model, we tailor an observation noise scalar prediction independently to capture how much the noise is contained in feature supervision during training, which is termed uncertainty-aware learning. Moving one step forward, as can be observed the distribution in terms of Stable Diffusion shows a bit shift away from the center scale, and usually carries broader value ranges and heavy-tailed (“spike”) values in the projected feature embeddings. A de-mean operation is then adopted to re-center the feature scales, being able to reduce the impact from anomaly points while still allowing points with small scale to guide the 3D model training.
@@
  • We analyze the heterogeneous yet complementary feature embeddings from multiple 2D models and incorporate both semantic- and geometric-aware knowledge into one single 3D model.
  • We further propose a deterministic uncertainty estimation to enable the 3D model predict independent observation scalar to capture the noise and resolve the heterogeneity from various feature supervisions.
@@ 2 Related Works
Open-Vocabulary (OV) 3D scene understanding advances over the previous large corpus of close-set approaches [8, 51, 70, 2, 98, 50, 87], allowing robust zero-shot reasoning and alleviating the need for annotations. Recent advances in Visual-Language Models (VLMs) [66, 34] have driven OV models towards remarkable levels of robustness with numerous emerging approaches tackling OV in image semantic segmentation [20, 5, 49, 86], object detection [104, 3], and recently universal segmentation [93]. Differently, OV for 3D scene understanding (OV3D) is limited in the data availability for training a purely fundamental 3D VLM. Alternatively, the community achieves OV3D by distilling zero-shot knowledge from recent VLMs [66, 5] and by mapping point cloud features to a queryable CLIP space. In 3D semantic segmentation, ConceptFusion [33] fuses VLM representations from multiple views into 3D points. Some methods [79] extends OV to 3D instance semantic segmentation based on CLIP [66] or SAM [38] to align the 3D space with language space while forcing instance-mask constraint. However, the recent OV3D methods heavily rely on the zero-shot knowledge of the 2D VLMs without investigating the reliability of the projected 2D feature representation. In this work, we move forward and explore how to aggregate knowledge from various 2D foundation models.
@@
Knowledge distillation (KD) [27, 62] aims at training compact student models with the supervision from more powerful and larger teacher models [45]. Introduced by the first work [27], the student model is trained to mimic the prediction behavior of the teacher model and has been extensively explored in subsequent works [69, 91, 1, 97, 47, 99], which has been applied successfully in a wide range of tasks going from supervised-training [56, 85, 81], to network compression [75, 7, 65] and to domain adaptation [100, 28, 26]. Recently, following the explosion of Visual-Language Models like CLIP [66] and ALIGN [34], KD has been introduced to efficiently transfer knowledge between different modalities [45, 52, 23], and recently to bridge the gap between the text and 3D point cloud modalities [9, 90, 32, 96, 64, 15, 88]. Recently, AM-RADIO [67] describes a general methodology for distilling multiple distinct foundation models into one, but still focus on only 2D domain. We primarily focus on studying and tackling the ambiguity of the distilled representations between image and point cloud modalities.
@@
Uncertainty estimation has been widely investigated in various tasks [35, 29, 40, 42, 60, 24, 57, 36] which is capable of addressing the problem of quantifying the uncertainty of predictions by model. Uncertainty Estimation can be broadly classified into: (i) aleatoric estimation [35, 83, 59] that usually dues to the underlying uncertainty in the measurement which utilizes the extra network to be trained from scratch to approximate a heteroscedastic distribution by maximizing the likelihood of the system, and (ii) epistemic estimation [22, 40, 19] that induces the uncertainty by the model parameters in low-data regimes as parameter estimation becomes noisy, respectively. In 3D point clouds, uncertainty estimation finds applications in incremental learning [89], semantic segmentation [12], and domain adaptation [73, 84, 74]. Unlike previous works, we seek a deterministic uncertainty estimation for supervising 3D model training under uncertainty awareness for the ambiguity between 2D image and 3D point modalities.
@@ 3 Methodology
In this section, we first describe the preliminary open-vocabulary 3D scene understanding task in Sec. 3.1. Then we elucidate inconsistent results across multi-view posed images to demonstrate the necessity of uncertainty-aware training to alleviate such issues and embrace various 2D foundation models in Sec. 3.2. Our Cross-Modal and Uncertainty-Aware Agglomeration (CUA-O3D) method will be depicted in Sec. 3.3, including Distillation agglomeration and Deterministic uncertainty estimation.
@@ 3.1 Open-Vocabulary 3D Scene Understanding
In standard 3D semantic segmentation, the training set $\mathcal{T}_{train}$ includes point clouds and dense point-level annotations. Each point cloud $\mathcal{X}={p_{i}\in\mathbb{R}^{3},i\in[0,N-1]}$, consists of $N$ points $p_{i}$ with corresponding point-level annotations $\mathcal{Y}$. These annotations are assigned from a predefined set of class indices $\mathcal{K}=[1,\dots,K]$, where each index corresponds to a specific class name in the vocabulary $\mathcal{V}=[v_{1},\dots,v_{k}]$. Given $\mathcal{T}_{train}$ and $\mathcal{K}$, the objective is to learn a deep neural network correctly assigning a label from $\mathcal{K}$ to each point $p_{i}\in\mathcal{X}$. In contrast, open-vocabulary 3D semantic segmentation (OV3D) aims to segment $\mathcal{X}$ using an arbitrary vocabulary $\mathcal{V}$. Recent approaches achieve this with a pre-trained VLM [64, 79]. The VLM provides this common embedding space through two distinct encoders, namely, a vision encoder $f^{2D}:\mathcal{I}\rightarrow\mathbf{Z}$ and a text encoder $f^{txt}:\mathcal{V}\rightarrow\mathbf{Z}$. The common practice is to train a 3D vision encoder $\theta_{3D}$ to align to $f^{2D}$ embeddings, bridging the modality gap and defining $\theta_{3D}:\mathcal{X}\rightarrow\mathbf{Z}$. After training, $f^{txt}$ encodes $\mathcal{V}$, and class predictions are computed via similarity matching between 3D point cloud and textual vocabulary embeddings.
@@ 3.2 Preliminary Observation
We first conduct a simple qualitative study to analyze how embedding ambiguity affects $f^{2D}$ predictions before and after projection to the point cloud space based on commonly used Lseg [5]. We analyze the consistency of the predictions for the same object appearing in multi-view images. Given a pre-trained vision-language model $f^{2D}$ and multi-view images $\mathcal{I}$, we query $f^{2D}$ with known vocabularies $\mathcal{V}$ and report the qualitative results over multiple ScanNetV2 [13] view-images in Fig. 2. We notice that predictions are consistent for the classes wall, floor, and door while showing inconsistency over the class cabinet. After projection to the 3D point cloud space (left), the projected prediction inherits this ambiguity, resulting in further detrimental 3D model distillation. Regarding the incorporation of various 2D foundation models shown in Fig. 1, how to harmonize the heterogeneous characteristics also matters. This simple study highlights the need for a reliable uncertainty measure capturing embedding ambiguity that we hope to shed new light on the future works on this task.
@@ 3.3 Cross-Modal Agglomeration
Distillation agglomeration. The distillation phase aims to align the point embeddings $F^{3D}$ to image embeddings $F^{2D}$ obtained from each frozen pre-trained visual encoder $f^{2D}_{i}$. This is a common practice in recent OV3D approaches [79], which leads to a shared embedding space between image, text, and point cloud modalities. Given a point cloud $\mathcal{X}$ paired with multi-view posed images $\mathcal{I}$, we learn 3D sparse features $F^{3D}$ by employing the well-established MinkowskiNet [11] as a 3D sparse convolutional encoder $\theta_{3D}$. The encoder $\theta_{3D}$ outputs a sparse set of point-wise feature vectors $F^{3D}=\theta_{3D}(\mathcal{X})$, where each feature is associated with an input point $p_{i}\in\mathcal{X}$. Similarly, we distill the multi-view image features from $\mathcal{I}$ using each frozen vision encoder $f^{2D}_{i}$ separately, where $\{f^{2D}_{i}\in f^{2D}|f^{2D}_{Lseg},f^{2D}_{DINOv2},f^{2D}_{SD}\}$. The output of $f^{2D}$ is a set of feature vectors $F^{2D}_{i}=f^{2D}_{i}(\mathcal{I})$ where each feature vector in $F^{2D}$ is associated to an input pixel $u$. After projection in the homogeneous coordinates space, we redefine $F^{2D}_{i}$ as a set of point-wise image features, corresponding to each $f^{2D}_{i}$. We define the matches between each point $p$ and pixel $u$ by using the corresponding homogeneous coordinates $\tilde{p}$ and $\tilde{u}$, respectively. Once matches are established, we enforce the alignment between the predicted $F^{3D}_{i}$ and $F^{2D}_{i}$. Following [64] and training $\theta_{3D}$ to minimize the distillation loss $\mathcal{L}_{distill}$, our independent distillation losses are defined as,
@@
which corresponds to minimizing the distance between each $F^{3D}_{i}$ and $F^{2D}_{i}$, and leading to final distillation loss as,
@@
Deterministic uncertainty estimation. As we analyzed before that different 2D foundation models encapsulate unique characteristics and one single 2D model induces inherent inconsistency from multi-view posed image which necessitates appropriate measures to tackle. We then propose a simple yet effective deterministic uncertainty-aware observation scalar prediction to quantify embedding ambiguity within each cross-modal distillation, that learns the adaptive weights of various 2D feature supervisions under the cross-modal training.
Specifically, we devise the 3D model $\theta_{3D}$ with three independent noise scalar predictions $\sigma_{i}$  w.r.t each 2D model $f^{2D}_{i}$. The output from the probabilistic model with weight $\mathrm{W}$ being analogous to the regression task can be modeled as gaussian likelihood as:
$p(y|f^{\mathrm{W}}(x))=N(f^{\mathrm{W}}(x),\sigma^{2}),$ (6)
@@
$p(y_{1},...,y_{K}|f^{\mathrm{W}}(x))=\prod\limits_{i=1}^{K}p(y_{i}|f^{\mathrm{W}}(x)),$ (7)
where index $i$ corresponds to the mapping for each 2D model, $\{f^{2D}_{Lseg},f^{2D}_{DINOv2},f^{2D}_{SD}\}$. Based on the maximum likelihood modeling and taking the regression task as an example, the log-likelihood of the model can then be optimized,
$\log p(y|f^{\mathrm{W}}(x))\propto-\frac{1}{2\sigma^{2}}||y-f^{\mathrm{W}}(x)||^{2}-\log\sigma,$ (8)
where $\sigma$ denotes the model’s observation noise scalar, being responsible for capturing the inherent uncertainty within 2D feature supervisions, which is mainly caused by heterogeneous and noisy feature embeddings from various 2D foundation models. Assuming that we have three outputs corresponding to Lseg, DINOv2, and Stable Diffusion, each following gaussian-like distribution, we then have:
$\begin{split}p(y_{1},y_{2},y_{3}|f^{\mathrm{W}}(x))&=\prod\limits_{i=1}^{3}p(y_{i}|f^{\mathrm{W}}(x))\\ &=\prod\limits_{i=1}^{3}N(y_{i};f^{\mathrm{W}}(x),\sigma_{i}^{2}),\end{split}$ (9)
and then, we formulate our training objective $\mathcal{L}_{distill}(\mathrm{W},\sigma_{1},\sigma_{2},\sigma_{3})$ for multiple mappings:
$\vspace{-0.2cm}\begin{split}&\mathcal{L}_{distill}=-\log p(y_{1},y_{2},y_{3}|f^{\mathrm{W}}(x))\\ &\propto\frac{1}{2\sigma_{1}^{2}}\mathcal{L}_{cos\_lseg}+\frac{1}{2\sigma_{2}^{2}}\mathcal{L}_{l1}+\frac{1}{2\sigma_{3}^{2}}\mathcal{L}_{cos\_sd}+\log\sigma_{1}\sigma_{2}\sigma_{3},\end{split}$ (10)
@@
To the best of our knowledge, this is the first work to explore deterministic uncertainty-aware modeling for agglomerating multiple 2D foundation models into a single unified 3D model, aiming toward the development of a potential foundational 3D model. In the last term of Eq. 10, each supervision signal from a 2D model contributes to learning adaptive weighting during training. Specifically, the uncertainty parameter $\sigma_{i}$, which characterizes the noise level associated with the $i$-th 2D model $f^{2D}_{i}$, dynamically adjusts the contribution of the corresponding loss term $\mathcal{L}_{i}$. As $\sigma_{i}$ increases—indicating higher uncertainty—the effective weight of $\mathcal{L}_{i}$ decreases, thereby down-weighting less reliable supervision. Meanwhile, each $\sigma_{i}$ is implicitly regularized to prevent excessive growth, ensuring that all supervision sources contribute meaningfully to the optimization. We visualize the evolution of $\sigma_{i}$ throughout training in the supplementary material. To further ensure stable optimization, we add a small constant (set to $1.0$) to each $\sigma_{i}$ to prevent the loss from becoming negative during training.
$\begin{split}\log\sigma_{i}\to\log(1.0+\sigma_{i}).\end{split}\vspace{-0.8cm}$ (11)